# Quantum Support Vector Regression for Robust Anomaly Detection

### v2

@@ Abstract
Anomaly Detection (AD) is critical in data analysis, particularly within the domain of IT security. In recent years, Machine Learning (ML) algorithms have emerged as a powerful tool for AD in large-scale data. In this study, we explore the potential of quantum ML approaches, specifically quantum kernel methods, for the application to robust AD. We build upon previous work on Quantum Support Vector Regression (QSVR) for semisupervised AD by conducting a comprehensive benchmark on IBM quantum hardware using eleven datasets. Our results demonstrate that QSVR achieves strong classification performance and even outperforms the noiseless simulation on two of these datasets. Moreover, we investigate the influence of – in the NISQ-era inevitable – quantum noise on the performance of the QSVR. Our findings reveal that the model exhibits robustness to depolarizing, phase damping, phase flip, and bit flip noise, while amplitude damping and miscalibration noise prove to be more disruptive. Finally, we explore the domain of Quantum Adversarial Machine Learning and demonstrate that QSVR is highly vulnerable to adversarial attacks and that noise does not improve the adversarial robustness of the model.
Index Terms:
benchmark, semisupervised learning, noise, hardware, adversarial attacks, robustness, quantum kernel methods, quantum machine learning
I Introduction
Quantum Machine Learning (QML) merges Quantum Computing (QC) and Machine Learning (ML) to exploit potential advantages of QC for ML. Recently, Quantum Kernel Methods (QKMs) have gained attention for their potential to replace many supervised QML models and their ability to surpass variational circuits, as evidenced by Schuld [supervised_qml_are_kernel_schuld2021]. Theoretical findings further prove that QKMs can potentially address classification problems intractable for classical ML, such as the discrete logarithm problem [quantum_kernel_advantage_discrete_log_2021].
@@ I Introduction
Anomaly Detection (AD) is crucial in the realm of IT security, as it identifies deviations from normal patterns in areas like network intrusion and fraud detection [Ruff2020]. However, it is important to note that ML models employed in security-sensitive contexts are susceptible to adversarial attacks, where small, carefully crafted perturbations in inputs can result in misclassification. (Quantum) Adversarial Machine Learning (QAML/AML) explores techniques for both generating these adversarial attacks and defending against them.
@@
Given the potential of QML to address problems challenging for classical methods, the application of QML to AD is a tempting progression. In 2023, Tscharke et al. [Tscharke] proposed a semisupervised AD approach based on the reconstruction loss of a Quantum Support Vector Regression (QSVR) equipped with a quantum kernel. The authors compared the performance of the QSVR against a Quantum Autoencoder (QAE), a classical Support Vector Regression (CSVR) with a Radial Basis Function (RBF) kernel, and a classical autoencoder (CAE), using ten real-world and one synthetic dataset. Their simulated QSVR demonstrated comparable performance to CSVR, with marginal superiority over the other models. However, the implementation of their QSVR on hardware was left open, a gap tackled by this paper.
In today’s Noisy Intermediate-Scale Quantum (NISQ) era, noise limits the application of QC for industrial tasks, underscoring the need for a comprehensive understanding of how noise affects quantum algorithms. Noise impacts QML models by affecting predictive performance and, in QAML, altering robustness against adversarial attacks [quantum_noise_against_adv_attacks_Du, noise_improve_adv_robustness_Huang].
The remainder of this work is structured as follows: The next subsection I-A offers a comprehensive review of research related to QKMs, the impact of noise on QML, and QAML. Our contributions are detailed in subsection I-B. The following Background (section II) provides a foundation for understanding QKMs, noisy QC, and adversarial attacks. Next, the Methods (section III) describe the implementation of the QSVR on hardware, the noise simulation, and the adversarial attacks generation. In Results and Discussion (section IV), we analyze the results of the hardware experiments and investigate the influence of noise and adversarial attacks on the model’s performance. Finally, Conclusion and Outlook (section V) highlights the key results of this work and provides future research directions.
I-A Related Work
@@
In 2019, Havlicek et al. [Havlicek2019] introduced a quantum Support Vector Machine (QSVM) for binary classification on a two-qubit NISQ device. Since then, QSVMs have been applied to many areas, including remote sensing image classification [Delilbasic2021], mental health treatment prediction [Ahmad2021], and breast cancer prediction [Mafu2021]. Kyriienko and Magnusson [quantum_one_class_SVM] extended this to unsupervised fraud detection with a simulated one-class QSVM, and Tscharke et al. [Tscharke] developed a QSVR for semisupervised AD in 2023. However, to the best of our knowledge, a QSVR for semisupervised AD has not yet been set up on hardware.
@@
Research has also explored the influence of noise in QML models [noise_influence_Diego, noise_influence_zhou, noise_influence_winderl], focusing on noise robustness [noise_robustness_Nguyen, noise_robustness_Yao] or beneficial use of noise [noise_beneficial_ju, noise_beneficial_winderl]. However, as far as we know, the influence of noise on a QSVR for semisupervised AD has not yet been evaluated.
@@
Finally, the link between quantum noise and QAML was established by Du et al. [quantum_noise_against_adv_attacks_Du] in 2021, who found that adding depolarization noise can increase adversarial robustness. Building on this, Huang and Zhang [noise_improve_adv_robustness_Huang] improved the adversarial robustness of Quantum Neural Networks by adding noise layers in 2023. To date, there have been no published results involving adversarial attacks on semisupervised QML models for AD.
I-B Contributions
@@
The goal of our work is to gain further insight into the potential of QSVR for semisupervised AD in the NISQ era, which we accomplish through these contributions:
@@
    We investigate how the model performs on hardware and report that the QSVR achieves good classification performance on a 27-qubit IBM device, even outperforming a noiseless simulation on two out of ten real-world datasets.
@@
    We show that the QSVR is largely robust to noise by training and evaluating over 500 noisy models. We further observe that amplitude damping and miscalibration have the most damaging effect on the model’s performance and that the artificial Toy dataset, constructed to be linearly separable, suffers the most from noise.
@@
    Finally, we investigate the robustness of the QSVR against adversarial attacks and find it highly vulnerable, with performance on real-world datasets dropping by up to an order of magnitude for a weak attack strength of $\varepsilon=0.01$. Introducing noise into the QSVR does not clearly improve the model’s adversarial robustness.
II Background
II-A Quantum Kernel Methods
A kernel is a positive, semidefinite function $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{C}$ on the input set $\mathcal{X}$. It uses a distance measure $\kappa(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ between two input vectors $\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\mathcal{X}$ to create a model that captures the properties of a data distribution. A feature map $\phi:\mathcal{X}\to\mathcal{F}$ maps input vectors $\boldsymbol{x}$ to a Hilbert or feature space $\mathcal{F}$. They are of great importance in ML, as they map input data in a higher-dimensional space with a well-defined metric. The feature map can be a nonlinear function that changes the relative position of the data points. As a result, the dataset can become easier to classify in feature space, and even linearly separable. We associate feature maps with kernels by defining a kernel via
@@ II-A Quantum Kernel Methods
where $\braket{\cdot,\cdot}_{\mathcal{F}}$ is the inner product defined on $\mathcal{F}$.
With the exponentially large Hilbert space of QCs, the use of QKMs for ML is close at hand. A quantum feature map $\psi:\boldsymbol{x}\to\ket{\phi(\boldsymbol{x})}$ is implemented via a feature-embedding circuit $U(\boldsymbol{x})$, which acts on a ground state |$\ket{0\dots 0}$ of a Hilbert space $\mathcal{F}$ as |$\ket{\phi(\boldsymbol{x})}=U(\boldsymbol{x})\ket{0\dots 0}$. The distance measure in the quantum kernel is the absolute square of the inner product of the quantum states. On hardware, this can be realized by the inversion test, where a sample $\boldsymbol{x}_{i}$ is encoded in the unitary $U$, followed by the adjoint $U^{\dagger}$ encoding the second sample $\boldsymbol{x}_{j}$ and measuring the probability of the all-zero state. Thus, the quantum kernel is defined as
$\kappa(x_{i},x_{j})$ $=\left|\braket{\phi(\boldsymbol{x}_{i})}{\phi(\boldsymbol{x}_{j})}\right|^{2}$
$=\left|\bra{0^{\otimes n}}U^{\dagger}(\boldsymbol{x}_{i})U(\boldsymbol{x}_{j})\ket{0^{\otimes n}}\right|^{2}$ (2)
@@
II-B Noisy Quantum Computing
Real quantum systems are never completely isolated from the environment; for example, an electron realizing a qubit will interact with other charged particles. Moreover, quantum computers are programmed by an external system and thus can never be a closed system [Nielsen, Suter_Noise].
In the current NISQ-era, noise significantly limits the performance of quantum algorithms, primarily through coherent and incoherent noise. Coherent noise arises from systematic, reversible errors that lead to predictable but undesired evolution of the states, often caused by imperfect calibrations or imprecise control signals [kaufmann2023_noise]. Coherent noise is an unitary evolution of the system, characterized by having only one operation element in the operator-sum representation introduced below [preskill_lectures_noise, Wallman_Noise].
Incoherent noise, on the other hand, is characterized by random, stochastic processes caused by insufficient isolation of the system from its environment. These uncontrolled interactions between system and environment lead to deviations between the desired and the actual evolution and to a loss of coherence in the system [kaufmann2023_noise, Suter_Noise].
Quantum noise can be modeled by a quantum channel, where the term ”channel” is drawn from classical information theory [Nielsen]. In the operator-sum representation, a channel is described by the map $\mathcal{E}$ with operation elements (or Kraus operators) $\{E_{i}\}$ mapping the density operator $\rho=\ket{\psi}\bra{\psi}$ | to another density operator $\mathcal{E}(\rho)$.
@@ II-B Noisy Quantum Computing
In the following, selected types of single-qubit noisy channels are described, and their operation elements $E_{i}$ are listed. For a more detailed explanation of the noisy quantum channels, we refer to [Nielsen, Wilde_2013, preskill_lectures_noise].
@@
    Amplitude Damping channel: describes the effect of energy loss, such as when an atom emits a photon. The channel acts on the quantum system $A$ and the environment $E$ as follows: if both $A$ and $E$ are in their ground state |$\ket{0}$, nothing happens. If $A$ is in the excited state |$\ket{1}_{A}$, a photon will be emitted with probability $p$, leading to the excitation of the environment and causing the transition |$\ket{0}_{E}\to\ket{1}_{E}$, while $A$ drops to the ground state, i.e. |$\ket{1}_{A}\to\ket{0}_{A}$. The evolution caused by the channel can be summarized as:
    $|0\rangle_{A}\otimes|0\rangle_{E}\mapsto|0\rangle_{A}\otimes|0\rangle_{E}$
    $|1\rangle_{A}\otimes|0\rangle_{E}\mapsto\sqrt{1-p}|1\rangle_{A}\otimes|0\rangle_{E}+\sqrt{p}|0\rangle_{A}\otimes|1\rangle_{E}$ (5)
@@
    Bitflip channel: flips the state of a qubit from |$\ket{0}$ to |$\ket{1}$ and vice versa with probability $p$. The operators are:
@@
    Depolarizing channel: the qubit remains intact with probability $1-p$, while an error occurs with probability $p$. If an error occurs, the state is replaced by a uniform ensemble of the three states $X\ket{\psi},Y\ket{\psi},Z\ket{\psi}$. This is a symmetric decoherence channel defined by operation elements:
@@
    Phaseflip channel: applies a phase of $-1$ to the |$\ket{1}$-state with probability $p$ and leaves the |$\ket{0}$-state unchanged. It has the operation elements:
@@ II-C Adversarial Attacks
III Methods
The experiments performed in this work are threefold. First, we benchmarked the QSVR for semi-supervised AD on hardware. Second, we evaluated the influence of noise on the classification performance of the QSVR, and third, we investigated the influence of noise on the adversarial robustness of the QSVR. An overview of the datasets used in the experiments is given in Table VI in Appendix A in the appendix. The datasets were reduced to five dimensions using Principal Components Analysis. For a more detailed description of the model, the datasets, and the preprocessing techniques, we refer to [Tscharke].
III-A Quantum Support Vector Regression Model
The QSVR for semi-supervised AD is described in detail in [Tscharke] , and the kernel circuit is displayed in Fig. 1. The first data point $\boldsymbol{x}_{i}$ is encoded by the unitary $U$, followed by its inverse $U^{\dagger}$ encoding the second data point $\boldsymbol{x}_{j}$. The unitary $U$ consists of a layer of RZ gates and a layer of RX gates, followed by a layer of IsingZZ [^1] https://docs.pennylane.ai/en/stable/code/api/pennylane.IsingZZ.html gates to create entanglement. Each of the single-qubit gates encodes one feature, while the parameter of the IsingZZ gate is a product of two features. The kernel entry $K_{ij}=K_{ji}$ is obtained by measuring the probability of the all-zero state after applying both unitaries.
@@ III-B Hardware Experiments
III-B1 Device Specifications
The experiments were performed on the IBM System One in Ehningen, Germany, in January 2024. The system is a 27-qubit superconducting quantum computer with a quantum volume of 64. The QSVR was implemented using qiskit [qiskit] with default error mitigation techniques. Further specifications of the system are listed in table I.
TABLE I: Specification of the IBM System One Ehningen.
@@ III-B1 Device Specifications
Name IBM Quantum System One at Ehningen
@@
Operation time of 2 qubit gate $\approx 300$ns for CNOT
III-B2 Reference Models
@@
III-C Generation of Noise
The influence of noise on the QSVR for semisupervised AD was evaluated by applying six noise channels with different strengths to the quantum circuit that computes the kernel. The seven noise probabilities used in the experiments are $p\in\{0.01,0.05,0.1,0.2,0.3,0.4,0.5\}$ and the five noise channels are [bitflip, phaseflip, depolarizing, phase damping, amplitude damping]. This leads to a total of $7\cdot 5=35$ models per dataset. For the miscalibration channel, the noise probability $p$ is the overrotation in radians in 20 linear steps between 0 and $2\pi$. For the DoH dataset subject to adversarial attacks of strength $\varepsilon=0.1$, additional evaluations were performed in the region close to $p=\pi$. The noisy QSVR was simulated using PennyLane [pennylane]. We used a training set of size 100 from the normal class and a test set of size 100 with a balanced class ratio.
III-D Generation of Adversarial Attacks
We created 100 adversarial samples of the test set (50 from each class) using PGD with the parameters listed in Table II. The attacks targeted the noiseless models and were then applied to the noisy models. For adversarial training, we create adversarial samples of the training set and train the model using the adversarial training set of size 100.
TABLE II: Overview of the parameters used in the PGD attacks.

### v3

# Anomaly Detection with Quantum SVR in the NISQ Era: Limits of Robustness to Noise and Adversarial Attacks

@@ Abstract
Anomaly Detection (AD) is critical in data analysis, particularly within the domain of IT security. In this study, we explore the potential of Quantum Machine Learning for application to AD with special focus on the robustness to noise and adversarial attacks. We build upon previous work on Quantum Support Vector Regression (QSVR) for semisupervised AD by conducting a comprehensive benchmark on IBM quantum hardware using eleven datasets. Our results demonstrate that QSVR achieves strong classification performance and even outperforms the noiseless simulation on two of these datasets. Moreover, we investigate the influence of – in the NISQ-era inevitable – quantum noise on the performance of the QSVR. Our findings reveal that the model exhibits robustness to depolarizing, phase damping, phase flip, and bit flip noise, while amplitude damping and miscalibration noise prove to be more disruptive. Finally, we explore the domain of Quantum Adversarial Machine Learning by demonstrating that QSVR is highly vulnerable to adversarial attacks, with neither quantum noise nor adversarial training improving the model’s robustness against such attacks.
1 INTRODUCTION
Quantum Kernel Methods (QKMs) have gained attention in Quantum Machine Learning (QML) as they can replace many supervised QML models [supervised_qml_are_kernel_schuld2021]. Theoretically, QKMs offer a promising route to tackle classification problems that are intractable for classical methods. In particular, Support Vector Machines (SVMs) with quantum kernels (QSVM) have been shown to achieve exponential speedups on certain tasks, such as classification based on the discrete logarithm problem [quantum_kernel_advantage_discrete_log_2021] or the k-Forrelation problem [J_ger_2023].
@@ 1 INTRODUCTION
Anomaly Detection (AD) is crucial in IT security, where it detects deviations from normal behavior in domains like intrusion and fraud detection [Ruff2020]. However, ML models are vulnerable to adversarial attacks [adv_attacks_szegedy] – small, crafted perturbations that cause misclassification – studied in Adversarial Machine Learning (AML).
@@
Given the potential of QML to address problems challenging for classical methods, the application of QML – and especially QKMs – to AD is a tempting progression. [Tscharke] proposed a semisupervised AD approach based on the reconstruction loss of a Quantum Support Vector Regression (QSVR) equipped with a quantum kernel. Their choice of QSVR was motivated by its close connection to the QSVM, which has demonstrated theoretical advantages over classical methods in specific problem settings. Moreover, the regression formulation offers a natural way to produce a continuous anomaly score, providing a quantitative measure of how strongly an input deviates from expected normal behavior. The authors compared the performance of the QSVR against classical and quantum baselines on eleven datasets. Their simulated QSVR demonstrated comparable performance to a classical SVR, with marginal superiority over the other models. However, their work did not address implementation on quantum hardware, nor did it investigate the adversarial robustness of the QSVR, two gaps this paper tackles to assess the model’s reliability in security-critical applications.
In the current Noisy Intermediate-Scale Quantum (NISQ)-era, hardware limitations such as decoherence and gate errors pose significant challenges for deploying quantum algorithms in practical settings. These noise effects not only degrade the performance of QML models but also influence their susceptibility to adversarial attacks. As shown in studies [quantum_noise_against_adv_attacks_Du, noise_improve_adv_robustness_Huang], noise can both undermine and, paradoxically, enhance model robustness depending on its nature and interaction dynamics. This dual role underscores the importance of investigating how noise impacts both the classification performance and adversarial resilience of quantum models like the QSVR, two aspects that this paper examines.
The remainder of this work is structured as follows. Section 1.1 and Section 2 review related work and provide the necessary preliminaries. Section 3 describes the experimental setup, Section 4 presents and discusses the results, and Section 5 concludes with future directions.
1.1 Related Work
@@
[Havlicek2019] introduced a QSVM for binary classification on two qubits of a five-qubit NISQ device. Since then, QSVMs have been applied to many areas, including remote sensing image classification [Delilbasic2021], mental health treatment prediction [Ahmad2021], and breast cancer prediction [Mafu2021]. [quantum_one_class_SVM] extended this to unsupervised fraud detection with a simulated one-class QSVM, and [Tscharke] developed a QSVR for semisupervised AD in 2023. However, a QSVR for semisupervised AD has not yet been set up on hardware.
@@
Research has also explored the influence of noise on QML models [noise_influence_Diego, noise_influence_winderl], focusing on noise robustness [noise_robustness_Nguyen, noise_robustness_Yao] or beneficial use of noise [noise_beneficial_ju, noise_beneficial_winderl]. However, the influence of noise on a QSVR for semisupervised AD has not yet been evaluated.
@@
Finally, the link between quantum noise and QAML was established by [quantum_noise_against_adv_attacks_Du], who found that adding depolarization noise can increase adversarial robustness. Building on this, [noise_improve_adv_robustness_Huang] improved the adversarial robustness of Quantum Neural Networks by adding noise layers. To date, there have been no published results involving adversarial attacks on semisupervised QSVR for AD.
1.2 Contributions
@@
The goal of our work is to gain further insight into the potential of QSVR for semisupervised AD in the NISQ-era, which we accomplish through these contributions:
@@
    We investigate how the model performs on hardware and report that the QSVR achieves good classification performance on five qubits of a 27-qubit IBM device, even outperforming a noiseless simulation on two out of eleven datasets.
@@
    We show that the QSVR is largely robust to noise by simulating over 500 noisy models. We further observe that amplitude damping and miscalibration have the most damaging effect on the model’s performance and that the artificial Toy dataset, constructed to be linearly separable, suffers the most from noise.
@@
    Finally, we investigate the robustness of the QSVR against adversarial attacks and find it highly vulnerable, with performance on real-world datasets dropping by up to an order of magnitude for a weak attack strength of $\varepsilon=0.01$. Introducing noise into the QSVR or carrying out adversarial training does not clearly improve the model’s adversarial robustness.
2 BACKGROUND
2.1 Quantum Kernel Methods
A kernel is a positive, semidefinite function $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ on the input set $\mathcal{X}$. It uses a similarity measure $\kappa(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ between two input vectors $\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\mathcal{X}$ to create a model that captures the properties of a data distribution. A feature map $\phi:\mathcal{X}\to\mathcal{F}$ maps input vectors $\boldsymbol{x}$ to a Hilbert or feature space $\mathcal{F}$. They are of great importance in ML, as they map input data in a higher-dimensional space. The feature map can be a nonlinear function that changes the relative position of the data points. As a result, the dataset can become easier to classify in feature space, and even linearly separable. We associate feature maps with kernels by defining a kernel via
@@ 2.1 Quantum Kernel Methods
[…]
$\kappa(x_{i},x_{j})$ $=\left|\braket{\phi(\boldsymbol{x}_{i})|\phi(\boldsymbol{x}_{j})}\right|^{2}$
[…]
@@
2.2 Noisy Quantum Computing
Real quantum systems are never completely isolated from the environment; for example, an electron realizing a qubit will interact with other charged particles. Moreover, quantum computers are programmed by an external system and thus can never be a fully closed system [Nielsen, Suter_Noise].
In the current NISQ-era, noise significantly limits the performance of quantum algorithms, primarily through coherent and incoherent noise. Coherent noise arises from systematic, reversible errors that lead to predictable but undesired evolution of the states, often caused by imperfect calibrations or imprecise control signals [kaufmann2023_noise]. It is a unitary evolution of the system, described by only one operation element in the operator-sum representation introduced below [preskill_lectures_noise, Wallman_Noise].
Incoherent noise, on the other hand, rises from random, stochastic processes caused by insufficient isolation of the system from its environment. These uncontrolled interactions between system and environment lead to deviations between the desired and the actual evolution of the system and to a loss of coherence [kaufmann2023_noise, Suter_Noise].
Quantum noise can be modeled by a quantum channel, where the term ”channel” is drawn from classical information theory [Nielsen]. In the operator-sum representation, a channel is described by the map $\mathcal{E}$ with operation elements (or Kraus operators) $\{E_{i}\}$ mapping the density operator $|\rho=\ket{\psi}\bra{\psi}$ to another density operator $\mathcal{E}(\rho)$.
@@ 2.2 Noisy Quantum Computing
In the following, selected types of single-qubit noisy channels are described, and their operation elements $E_{i}$ are listed. For a more detailed explanation of the noisy quantum channels, we refer to [Nielsen, preskill_lectures_noise].
@@
[…]
    $|0\rangle_{A}\otimes|0\rangle_{E}\mapsto|0\rangle_{A}\otimes|0\rangle_{E}$ (5)
    $|1\rangle_{A}\otimes|0\rangle_{E}\mapsto\sqrt{1-p}|1\rangle_{A}\otimes|0\rangle_{E}+\sqrt{p}|0\rangle_{A}\otimes|1\rangle_{E}$
@@
[…]
@@
[…]
@@
[…]
@@ 2.3 Adversarial Attacks
3 METHODS
The experiments performed in this work are threefold. First, we benchmarked the QSVR for semi-supervised AD on hardware. Second, we evaluated the influence of noise on the classification performance of the QSVR, and third, we investigated the influence of noise and adversarial retraining on the adversarial robustness of the QSVR. An overview of the datasets used in the experiments is given in Table 6 in the appendix. The datasets were reduced to five dimensions using Principal Components Analysis. For a more detailed description of the model, the datasets, and the preprocessing techniques, we refer to [Tscharke].
3.1 Quantum Support Vector Regression Model
The QSVR architecture for semi-supervised AD was originally introduced in [Tscharke], and the corresponding kernel circuit is illustrated in Figure 1. To compute a kernel entry, the data point $\boldsymbol{x}_{i}$ is first encoded by the unitary $U$, followed by the application of its inverse $U^{\dagger}$ embedding a second data point $\boldsymbol{x}_{j}$. Our $U$ consists of a layer of RZ and RX gates, followed by IsingZZ [^1] https://docs.pennylane.ai/en/stable/code/api/pennylane.IsingZZ.html gates that generate entanglement. Each single-qubit rotation encodes one feature, while the parameter of each IsingZZ gate is given by the product of two features. The kernel value $K_{ij}=K_{ji}$ is obtained by measuring the probability of the all-zero state after execution of $U^{\dagger}U$.
@@ 3.2 Hardware Experiments
3.2.1 Device Specifications
The experiments were performed on the IBM System One in Ehningen, Germany. The system is a 27-qubit superconducting quantum computer with a quantum volume of 64. The QSVR was implemented using qiskit [qiskit] with default error mitigation techniques. Further specifications of the system are listed in table 1.
Table 1: Specification of the IBM System One Ehningen.
@@ 3.2.1 Device Specifications
Name IBM Quantum Sys-
tem One at Ehningen
@@
Operation time of CNOT $\approx 300$ns
3.2.2 Reference Models
@@
3.3 Generation of Noise
The influence of noise on the QSVR for semisupervised AD was evaluated by applying six noise channels with different strengths to the quantum circuit that computes the kernel. The seven noise probabilities used in the experiments are $p\in\{0.01,0.05,0.1,0.2,0.3,0.4,0.5\}$ and the five incoherent noise channels are [bit flip, phase flip, depolarizing, phase damping, amplitude damping]. This leads to a total of $7\cdot 5=35$ models per dataset. For the miscalibration channel, the noise probability $p$ is the overrotation in radians in 20 linear steps between 0 and $2\pi$. For the DoH dataset subject to adversarial attacks of strength $\varepsilon=0.1$, additional evaluations were performed in the region close to $p=\pi$. The noisy QSVR was simulated using PennyLane [pennylane]. We used a training set of size 100 from the normal class and a test set of size 100 with a balanced class ratio.
3.4 Generation of Adversarial Attacks
We created 100 adversarial samples of the test set (50 from each class) using PGD with the parameters listed in Table 2. The attacks targeted the noiseless models and were then applied to the noisy models. For adversarial training, we create adversarial samples of the training set and train the model using the adversarial training set of size 100.
Table 2: Overview of the parameters used in the PGD attacks.